OpenStudy (hesan):

lim_{x rightarrow infty}(1+1/x)^lnx Solve this....

OpenStudy (anonymous):

\[\lim_{x\to\infty}\left(1+\frac{1}{x}\right)^{\ln x}=\lim_{x\to\infty}1^{\ln x}=1\]

OpenStudy (hesan):

I have another question which based on Probability